# Rejection via Learning Density Ratios

### v1

@@ 4 Learning Idealized Distributions
Suppose $\alpha>1$ and $\lambda>0$, then Eq. 14 simplifies to,
$\rho^{\alpha}_{\lambda}(x)=\left[\left(\frac{\alpha-1}{2}\cdot\left(b-\frac{L^{\prime}(x)}{\lambda}\right)\right)^{\frac{2}{\alpha-1}}\right]_{+},$ (15)
where we take non-integer powers of negative values as 0.
@@
Both Corollaries 4.6 and 4.7 can provide a unique rejector policy than the KL-divergence variant. Corollary 4.7 can provide a similar effect when $a(x)\neq 0$ for all $x$. Nevertheless, having to determine which inputs $a(x)\neq 0$ and solving these values are difficult in practice. As such, we focus on $\alpha>0$. If there are values of $L^{\prime}(x)$ with high risk, the $\max$ will flatten these inputs to 00 in Corollary 4.6. However, if the original model $h$ performs well and $L^{\prime}(x)$ is relatively small for all $x$, then it is possible that the $\max$ is not utilized. In such a case, the $\alpha$-divergence rejectors can end up being similar — this follows the fact that (as $\varphi_{\alpha}^{\prime\prime}(1)>0$) locally all $\alpha$-divergences will be similar to the $\chi^{2}$ / $(\alpha=3)$-divergence [46, Theorem 7.20]. This ultimately results in $\alpha$-divergences being similar to, e.g., Chow’s rule when $h_{y}(x)\approx\Pr(\mathsf{Y}=y\mid\mathsf{X}=x)$ in classification via Theorem 2.1.
@@
Practical Rejection
In the following, we consider practical concerns for utilizing rejector for the KL-or $(\alpha>1)$-divergence case, Eqs. 13 and 15. To do so, we need to estimate the loss $L^{\prime}(x)=\mathbb{E}_{\mathrm{P}_{\rm y\mid\rm x}}\left[\ell(\mathsf{Y},h(x))\right]$ and rejector normalizer $Z$ or $b$.
Loss:   The former is tricky, we require an estimate to evaluate $\mathbb{E}_{\mathrm{P}_{\rm y\mid\rm x}}\left[\ell(\mathsf{Y},h(x))\right]$ over any possible $x\in\mathcal{X}$ to allow us to make a rejection decision. Implicitly, this requires us to have a high quality estimate of $\mathrm{P}_{\rm y\mid\rm x}$. In a general learning setting, this can be difficult to obtain — in fact it is just the overall objective that we are trying to learning, i.e., predicting a target $y$ given an input $x$. However, in the case of CPE (classification), it is not unreasonable to obtain an calibrated estimate of $\mathrm{P}_{\rm y\mid\rm x}$ via the classifier $h\colon\mathcal{X}\rightarrow\mathcal{D}(\mathcal{Y})$ [45]. In Section 5, we utilize temperature scaling to calibrate the neural networks we learn to provide such an estimate [29]. Hence, we set $L^{\prime}(x)=\mathbb{E}_{{\mathsf{Y}\sim h(x)}}\left[\ell(\mathsf{Y},h(x))\right]$. For proper CPE loss functions, $\mathbb{E}_{{\mathsf{Y}\sim h(x)}}\left[\ell(\mathsf{Y},h(x))\right]$ acts as a generalized entropy function. As such, the rejectors Eqs. 13 and 15 act as functions over said generalized entropy functions.
@@
${\mathbb{E}}_{\hat{\mathrm{P}}}\left[\left(\frac{\alpha-1}{2}\cdot\left(b-\frac{L^{\prime}(x)}{\lambda}\right)\right)^{\frac{2}{\alpha-1}}\right]_{+}-1=0.$
@@
Cut-off $\tau$:   In addition to learning the density ratio, an additional consideration is how to tune $\tau$ in the rejection decision, Definition 3.2. Given a fixed density estimator $\rho$, change $\tau$ amounts to changing the rate of rejection. We note that this problem is not limited to our density ratio rejectors, where approaches with rejectors $r\in\mathcal{R}$ via a (surrogate) minimization of Eq. 2 may be required multiple rounds of training with different rejection costs $c$ to find an accept rate of rejection. In our case, we have a fixed $\rho$ which allows easy tuning of $\tau$ given a validation dataset, similar to other confidence based rejection approaches, e.g., tuning a threshold for the margin of a classifier [7].
@@ 5 Experiments
Figure 2: Accuracy vs coverage plots across all datasets and approaches, with 50 equidistant $\tau\in(0,1]$ and $c\in[0,0.5)$ values (sorted by coverage). The black horizontal line depicts base models trained without rejection. Missing approaches in the plots indicates that the model rejects more than 60% of test points or has accuracy below the base model. Shaded region indicates ±$\pm 1$ s.t.d. region.
In this section, we evaluate our distributional approach to rejection across a number of datasets. In particular, we consider the standard classification setting with an addition setting where uniform label noise is introduced [4, 27]. For our density ratio rejectors, we evaluate the KL-divergence based rejector (Corollary 4.1) and $(\alpha=3)$-based rejectors (Corollary 4.6) with 50 equidistant $\tau\in(0,1]$ values. For our tests, we fix $\lambda=1$. Throughout our evaluation, we assume that a neural network (NN) model without rejection is accessible for all (applicable) approaches and this model is calibrated via temperature scaling [29]. As such, we use the practical considerations discussed in Section 4 for our density ratio rejectors. For our density ratio rejectors, we utilize the log-loss.
@@
We consider 3 multiclass classification datasets. For tabular datasets, we consider the gas drift dataset [56] and the human activity recognition (HAR) dataset [5]. Each of these datasets consists of 6 class labels to predict. Furthermore, we utilize a two hidden layer multilayer perceptron NN model for these datasets. In addition we consider the MNIST image dataset [37], where we utilize a convolutional NN. These prediction models are trained utilizing the standard logistic / log-loss without rejection and then are calibrated. For each of these datasets, we utilize a clean and noisy variant. For the noisy variant, we flip the class labels of the train set with a rate of 25%. We note that the test set is clean in both cases. All evaluation uses 5-fold cross validation. All implementation are done using PyTorch and are trained using a p3.2xlarge instance on AWS.
We consider 4 different baseline for comparison. Each baseline is trained with 50 equidistant costs $c\in[0,0.5)$. One baseline used in comparison is Mozannar and Sontag [40]’s cross-entropy surrogate approach (DEFER) which reduces the $C$ multi-class classification problem into a $C+1$ multi-class classification problem, treating the rejection option as a separate class. A generalization of DEFER is also considered [13] where we change the surrogate loss of cross-entropy to generalized cross-entropy [61] (GCE). We also consider a cost-sensitive classification reduction (CSS) of the classification with rejection problem [14] utilizing the sigmoid loss function. The aforementioned 3 baselines all learn a model with rejection simultaneously, i.e., a pretrained model cannot be utilized. We also consider a two stage predictor-rejector approach (PredRej) which learns a rejector from a pretrained classifier [39].
@@ Results
Fig. 2 presents a summary plot of the acceptance coverage versus model accuracy after rejection on the test set. The plot is focused around the 60% to 100% coverage region, with an extend plot presented in Appendix P. Over all folds for MNIST our density ratio rejectors take approximately $\sim 1/2$ hour to fit. A single baseline (fixed $c$) takes upwards of 2 hour for a single fold. Overall, given that the underlying model is calibrated, we find that our KL-divergence and $\alpha=3$-divergence density ratio rejector are either competitive or superior to the baselines. The baselines seem to become more competitive as noise increases, see Appendix P for further details. Furthermore, one might notice that the aforementioned baselines do not or provide poor trade-offs for coverage values $>95\%$. Indeed, to achieve rejection with high coverage (without architecture tuning), approaches which ‘wrap’ a base classifier seem preferable, i.e., PredRej and our density ratios rejectors. It should be noted that if large models are allowed to be used for the rejector — as per the MNIST case — CSS, DEFER, and GCE can provide superior accuracy vs acceptance coverage trade-offs (Noisy MNIST).
Among the approaches which ‘wrap’ the base classifier $h$, we find that these approaches have higher variance ranges than the other approaches. In particular, the randomness of the base model potentially magnifies the randomness after rejection. The variance range of the base model tends to increase as the noise increases (additional ranges of noise for HAR and Gas Drift are presented in the Appendix). The influence on rejection is unsurprising as these ‘wrapping’ approaches predict via a composition of the original model (and hence inherits its randomness across folds). In general, our density ratio rejector outperforms PredRej. However, it should be noted that PredRej does not require a calibrated classifier. Among the density ratio rejectors, between KL and $\alpha=3$, the only variation is in coverage region that the $\tau\in(0,1]$ threshold covers. This follows from the fact that for similar distributions, $\varphi$-divergences act similarly [46, Theorem 7.20]. We find this pattern holds for other values of $\alpha$.
@@ 6 Limitations and Conclusions
We propose a new framework for rejection by learning idealized density ratios. Our proposed rejection framework links typically explored classification with rejection to generalized variational inference and distributionally robust optimization. It should be noted that although we have focused on classification, $L^{\prime}(\mathrm{Q})$ could in theory be replaced by any other loss functions. In this sense, one could adapt this approach to other learning tasks such as regression, discussed in Appendix N. Furthermore, although we have focused on $\varphi$-divergences, there are many alternative ways idealized distribution can be constructed, e.g., integral probability metrics [42, 9]. One limitation of our distributional way of rejecting is the reliance on approximating $\mathrm{P}(\mathsf{Y}\mid\mathsf{X})$ with model $h$. Alternatively, in future work, one may seek to approximate the density ratio $\rho$ by explicitly learning densities $\mathrm{Q}$ and $\mathrm{P}$ or via gradient based methods (for the latter, see Appendix O).
@@
$\hookrightarrow$ Appendix P: Additional Experimental Details Pg P
@@
$r^{\textsc{KL}}_{\tau}=\left\llbracket L^{\prime}(x)\leq-\lambda\left(\log Z+\log\tau\right)\right\rrbracket,$
@@
Appendix P Additional Experimental Details
P.I Training Settings
@@
P.II Extended Plots

### v2

@@ 4.2 Alpha-Divergences
Suppose $\alpha>1$ and $\lambda>0$, then Eq. 14 simplifies to, $\rho^{\alpha}_{\lambda}(x)=\left[\left(\frac{\alpha-1}{2}\cdot\left(b-\frac{L^{\prime}(x)}{\lambda}\right)\right)^{\frac{2}{\alpha-1}}\right]_{+},$ (15) where we take non-integer powers of negative values as 0.
@@
Both Corollaries 4.6 and 4.7 can provide a unique rejector policy than the KL-divergence variant. Corollary 4.7 can provide a similar effect when $a(x)\neq 0$ for all $x$. Nevertheless, having to determine which inputs $a(x)\neq 0$ and solving these values are difficult in practice. As such, we focus on $\alpha>0$. If there are values of $L^{\prime}(x)$ with high risk, the $\max$ will flatten these inputs to 00 in Corollary 4.6. However, if the original model $h$ performs well and $L^{\prime}(x)$ is relatively small for all $x$, then it is possible that the $\max$ is not utilized. In such a case, the $\alpha$-divergence rejectors can end up being similar — this follows the fact that (as $\varphi_{\alpha}^{\prime\prime}(1)>0$) locally all $\alpha$-divergences will be similar to the $\chi^{2}$ / $(\alpha=3)$-divergence [56, Theorem 7.20]. This ultimately results in $\alpha$-divergences being similar to, e.g., Chow’s rule when $h_{y}(x)\approx\Pr(\mathsf{Y}=y\mid\mathsf{X}=x)$ in classification via Theorem 2.1.
@@
4.3 Practical Rejection
We consider practical concerns for utilizing the KL-or ($\alpha$ > 1)-divergence case, Eqs. 13 and 15, for post-hoc rejection. To do so, we need to estimate the loss $L^{\prime}(x)$ and rejector normalizer $Z$ or $b$.
Loss:   The former is tricky, we require an estimate to evaluate $L^{\prime}(x)=\mathbb{E}_{\mathrm{P}_{\rm y\mid\rm x}}\left[\ell(\mathsf{Y},h(x))\right]$ over any possible $x\in\mathcal{X}$ to allow us to make a rejection decision. Implicitly, this requires us to have a high quality estimate of $\mathrm{P}_{\rm y\mid\rm x}$. In a general learning setting, this can be difficult to obtain — in fact it is just the overall objective that we are trying to learning, i.e., predicting a target $y$ given an input $x$. However, in the case of CPE (classification), it is not unreasonable to obtain an calibrated estimate of $\mathrm{P}_{\rm y\mid\rm x}$ via the classifier $h\colon\mathcal{X}\rightarrow\mathcal{D}(\mathcal{Y})$ [55]. In Section 5, we utilize temperature scaling to calibrate the neural networks we learn to provide such an estimate [32]. Hence, we set $L^{\prime}(x)=\mathbb{E}_{{\mathsf{Y}\sim h(x)}}\left[\ell(\mathsf{Y},h(x))\right]$. For proper CPE loss functions, $\mathbb{E}_{{\mathsf{Y}\sim h(x)}}\left[\ell(\mathsf{Y},h(x))\right]$ acts as a generalized entropy function. As such, the rejectors Eqs. 13 and 15 act as functions over said generalized entropy functions. It should be noted that simply considering the softmax outputs of neural networks for probability estimation have seen prior success [29, 39]. This is equivalent to taking the 0-1-loss function [15, 35, 28] and taking a plugin estimate of probabilities via the neural network’s output. The study of using plugin estimates have also been explored in the model cascade literature [40].
@@ 4.3 Practical Rejection
${\mathbb{E}}_{\hat{\mathrm{P}}}\left[\left(\frac{\alpha-1}{2}\cdot\left(b-\frac{L^{\prime}(x)}{\lambda}\right)\right)^{\frac{2}{\alpha-1}}\right]_{+}-1=0.$ (17)
@@
Threshold $\tau$:   In addition to learning the density ratio, an additional consideration is how to tune $\tau$ in the rejection decision, Definition 3.2. Given a fixed density estimator $\rho$, change $\tau$ amounts to changing the rate of rejection. We note that this problem is not limited to our density ratio rejectors, where approaches with rejectors $r\in\mathcal{R}$ via a (surrogate) minimization of Eq. 2 may be required multiple rounds of training with different rejection costs $c$ to find an accept rate of rejection. In our case, we have a fixed $\rho$ which allows easy tuning of $\tau$ given a validation dataset, similar to other confidence based rejection approaches, e.g., tuning a threshold for the margin of a classifier [7].
@@ 5 Experiments
Table 1: Summary of rejection methods over all baselines and datasets targeting 80% coverage. Each cell reports the “accuracy [coverage]” values, bold for most accurate, and s.t.d. reported in Appendix.
Base KL-Rej ($\alpha$=3)-Rej PredRej CSS DEFER GCE
Clean HAR 97.38 [100] 99.93 [81] 99.93 [82] 98.86 [85] 99.58 [81] 99.44 [80] 99.31 [82]
Gas Drift 94.10 [100] 99.16 [80] 99.16 [80] 98.12 [80] 98.68 [80] 98.06 [80] 97.62 [80]
MNIST 98.55 [100] 99.93 [87] 99.93 [88] 99.18 [74] 99.95 [83] 99.93 [80] 99.85 [80]
CIFAR-10 90.20 [100] 97.22 [80] 97.17 [81] 91.40 [74] 95.45 [81] 93.72 [81] 94.25 [81]
OrganMNIST 89.10 [100] 96.55 [80] 96.52 [80] 93.79 [82] 94.49 [80] 93.47 [80] 93.68 [80]
OctMNIST 91.93 [100] 97.08 [81] 97.18 [80] 93.43 [86] 95.40 [81] 94.66 [85] 94.91 [87]
Noisy (25%) HAR 96.51 [100] 98.56 [81] 98.56 [81] 97.22 [80] 97.82 [81] 97.78 [69] 98.85 [80]
Gas Drift 93.84 [100] 97.30 [80] 97.28 [80] 95.87 [82] 98.71 [80] 99.02 [77] 97.52 [75]
MNIST 97.88 [100] 99.89 [80] 99.89 [81] 98.00 [93] 99.94 [80] 99.93 [81] 99.95 [81]
CIFAR-10 85.31 [100] 92.25 [82] 92.50 [81] 85.84 [88] 89.58 [82] 90.93 [80] 92.22 [81]
OrganMNIST 89.10 [100] 95.86 [82] 96.29 [81] 93.40 [84] 96.74 [81] 94.67 [80] 94.48 [80]
OctMNIST 91.89 [100] 97.17 [80] 97.10 [81] 93.42 [83] 95.49 [81] 94.08 [89] 94.63 [78]
In this section, we evaluate our distributional approach to rejection across a number of datasets. In particular, we consider the standard classification setting with an addition setting where uniform label noise is introduced [4, 30]. For our density ratio rejectors, we evaluate the KL-divergence based rejector (Corollary 4.1) and ($\alpha$=3)-based rejectors (Corollary 4.6) with 50 equidistant $\tau\in(0,1]$ values. For our tests, we fix $\lambda=1$. Throughout our evaluation, we assume that a neural network (NN) model without rejection is accessible for all (applicable) approaches. For our density ratio rejectors, we utilize the log-loss, practical considerations in Section 4.3, and Algorithm 1.222Our rejector’s code public at: https://github.com/alexandersoen/density-ratio-rejection.
@@
We consider 6 multiclass classification datasets. For tabular datasets, we consider the gas drift dataset [68] and the human activity recognition (HAR) dataset [5]. Each of these datasets consists of 6 classes to predict. Furthermore, we utilize a two hidden layer MLP NN model for these datasets. We consider the MNIST image dataset [46] (10 classes), where we utilize a convolutional NN. Additionally, we consider 3 larger image datasets with ResNet-18 architecture [33]: CIFAR-10 [44] (10 classes); and OrgMNIST / OrganSMNIST (11 classes) and OctMNIST (4 classes) from the MedMNIST collection Yang et al. [69, 70]. These prediction models are trained utilizing the standard logistic / log-loss without rejection and then are calibrated via temperature scaling [32]. For each of these datasets, we utilize both clean and noisy variants. For the noisy variant, we flip the class labels of the train set with a rate of 25%. We note that the test set is clean in both cases. All evaluation uses 5-fold cross validation. All implementation use PyTorch and training was done on a p3.2xlarge AWS instance.
We consider 4 different baseline for comparison. Each is trained with 50 equidistant costs $c,\tau\in[0,0.5)$, except on OctMNIST which uses 10 equidistant costs (selecting $c,\tau$ discussed in Appendix P). One baseline used corresponds to a modification of Mozannar and Sontag [49]’s cross-entropy surrogate approach (DEFER) originally used for the learning to defer literature (see Charoenphakdee et al. [14, Appendix A.2]). This approach treats the rejection option as a separate class and solves a $|\mathcal{Y}|+1$ classification problem. A generalization of DEFER is considered which utilizes generalized cross-entropy [75] as its surrogate loss (GCE) [13]. We also consider a cost-sensitive classification reduction (CSS) of the classification with rejection problem [14] utilizing the sigmoid loss function. The aforementioned 3 baselines all learn a model with rejection simultaneously, i.e., a pretrained model cannot be utilized. We also consider a two stage predictor-rejector approach (PredRej) which learns a rejector from a pretrained classifier [48].
@@ Results
Table 1 presents a tabular summary of accuracy and coverage values when targeting 80% coverage values; and Fig. 2 presents a summary plot of the acceptance coverage versus model accuracy after rejection, focused around the 60% to 100% coverage region. This plot is limited to HAR, Gas Drift, and MNIST due to space limitations however the deferred datasets show curves where the density ratio rejector dominate with better accuracy and coverage in the plotted region, with corresponding extended plots for all datasets in Appendix Q . Over all folds for MNIST our density ratio rejectors take approximately $\approx 1/2$ hour to fit. A single baseline (fixed $c$) takes upwards of 2 hour for a single fold. Overall, given that the underlying model is calibrated, we find that our density ratio rejector are either competitive or superior to the baselines. One might notice that the aforementioned baselines do not or provide poor trade-offs for coverage values $>95\%$ (as per Fig. 2). Indeed, to achieve rejection with high coverage (without architecture tuning), approaches which ‘wrap’ a base classifier seem preferable, i.e., PredRej and our density ratios rejectors. Even at lower coverage targets (80%), Table 1 shows that density-ratio methods are comparable or superior in the more complex datasets of CIFAR-10 and the MedMNIST collection. If large models are allowed to be used for the rejector — as per the MNIST case — CSS, DEFER, and GCE can provide superior accuracy vs acceptance coverage trade-offs (noisy MNIST). However, this is not always true as per CIFAR-10 where all approaches are similarly effected by noise; or OctMNIST and OrgMNIST where approaches only slightly change with noise. The latter appears to be a consequence of label noise not effecting the Base classifier’s accuracy (using the larger ResNet-18 architecture), as per Table 1.
Among the approaches which ‘wrap’ the base classifier $h$, we find that these approaches have higher variance ranges than the other approaches. In particular, the randomness of the base model potentially magnifies the randomness after rejection. The variance range of the base model tends to increase as the noise increases (additional ranges of noise for HAR and Gas Drift are presented in the Appendix). The influence on rejection is unsurprising as these ‘wrapping’ approaches predict via a composition of the original model (and hence inherits its randomness across folds). In general, our density ratio rejector outperforms PredRej. However, it should be noted that PredRej does not require a calibrated classifier. Among the density ratio rejectors, between KL and $\alpha=3$, the only variation is in coverage region that the $\tau\in(0,1]$ threshold covers. This follows from the fact that for similar distributions, $\varphi$-divergences act similarly [56, Theorem 7.20]. We find this pattern holds for other values of $\alpha$.
Figure 2: Accuracy vs coverage plots across select datasets and all approaches, with 50 equidistant $\tau\in(0,1]$ and $c\in[0,0.5)$ values (sorted by coverage). The black horizontal line depicts base models trained without rejection. Missing approaches in the plots indicates that the model rejects more than 60% of test points or has accuracy below the base model. Shaded region indicates ±$\pm 1$ s.t.d. region.
@@ 6 Limitations and Conclusions
We propose a new framework for rejection by learning idealized density ratios. Our proposed rejection framework links typically explored classification with rejection to generalized variational inference and distributionally robust optimization. It should be noted that although we have focused on classification, $L^{\prime}(\mathrm{Q})$ could in theory be replaced by any other loss functions. In this sense, one could adapt this approach to other learning tasks such as regression, discussed in Appendix N. Furthermore, although we have focused on $\varphi$-divergences, there are many alternative ways idealized distribution can be constructed, e.g., integral probability metrics [51, 9]. One limitation of our distributional way of rejecting is the reliance on approximating $\mathrm{P}[\mathsf{Y}\mid\mathsf{X}]$ with model $h$. In future work, one may seek to approximate the density ratio $\rho$ by explicitly learning densities $\mathrm{Q}$ and $\mathrm{P}$ or via gradient based methods (for the latter, see Appendix O).
Acknowledgments and Disclosure of Funding
We thank the reviewers and the area chair for the various suggestions which improved the paper. A majority of the work was done whilst AS was interning at Amazon.
@@
$\hookrightarrow$ Appendix P: Finding the Best Rejection Cost Pg P
$\hookrightarrow$ Appendix Q: Additional Experimental Details Pg Q
@@
$r^{\textsc{KL}}_{\tau}(x)=\left\llbracket L^{\prime}(x)\geq-\lambda\left(\log Z+\log\tau\right)\right\rrbracket,$
@@
Appendix P Finding the Best Rejection Cost
Both the baselines and density ratio approaches evaluated in Section 5 require a tuning of hyperparameters $c,\tau$. Practically, we are more interested in the rejection rate $\mathrm{P}[r(\mathsf{X})=1]$ rather than the abstract choices of $c,\tau$. Hence, one often wishes to select $c,\tau$ according to a coverage constraint.
One of the simplest choices of selecting $c,\tau$ is to utilize a calibration set and approximate the rejection rate / coverage on this data. Our experiments in Section 5 mirrors this in Tables 1 and I by treating the final test set as the calibration data.
Other works in the literature utilize more sophisticated methods for deriving coverage constraints. Geifman and El-Yaniv [29] picks threshold values which provide a guarantee on a coverage-modified risk requirement. Another approach Pugnana and Ruggieri [57] does not require an additional calibration set and instead fits threshold values $\tau$ by exploiting stacking confidences and (stratified) cross-fitting. Although one could consider such approaches for all baselines, including those which utilize the cost-model of rejection via $c$, one will have to pay a price in retraining the learned rejector $r$ for each instance of the hyperparameter $c$ that is searched. As a result, approaches which can trade-off accuracy and coverage via a threshold variable $\tau$ (rather than an expensive retraining) are computationally preferable when exhaustively searching for a good $c,\tau$.
We leave more sophisticated methods for selecting $\tau$ for density ratio rejectors for future work.
Appendix Q Additional Experimental Details
Q.I Training Settings
@@
  • CIFAR-10 (CC BY 4.0): We utilize a ResNet-18 classifier. Random cropping and horizontal flipping data augmentation is utilized in training. We take a 256 batch size, 40 training epochs and a 0.0001 learning rate.
  • OrgMNIST (CC BY 4.0): We utilize a ResNet-18 classifier. We take a 256 batch size, 40 training epochs and a 0.0001 learning rate.
  • OctMNIST (CC BY 4.0): We utilize a ResNet-18 classifier. We take a 256 batch size, 40 training epochs and a 0.0001 learning rate.
@@
Q.II Extended Table and Plots
Table I presents an extended version of Table 1 over coverage targets of 80% and 90%. The standard deviation is additionally reported. One can see the observation from the main text are consistent with this extended table.
Table I: Extended Summary performance of rejection methods over all baselines and datasets targeting 80% and 90% coverage. Each cell reports the “accuracy (accuracy s.t.d.)[coverage (coverage s.t.d.)]” values.
Base KL-Rej ($\alpha$=3)-Rej PredRej CSS DEFER GCE
Target Coverage: 80% Clean HAR 97.38 (0.34) [100.00 (0.00)] 99.93 (0.06) [80.74 (0.84)] 99.93 (0.06) [81.82 (0.74)] 98.86 (0.44) [84.86 (8.45)] 99.58 (0.31) [81.39 (1.48)] 99.44 (0.19) [80.16 (1.54)] 99.31 (0.52) [81.78 (4.69)]
Gas Drift 94.10 (1.87) [100.00 (0.00)] 99.16 (0.62) [80.04 (0.78)] 99.16 (0.62) [80.24 (0.75)] 98.12 (1.03) [80.09 (6.03)] 98.68 (0.39) [80.33 (1.18)] 98.06 (0.97) [80.49 (1.09)] 97.62 (0.36) [80.43 (3.92)]
MNIST 98.55 (0.19) [100.00 (0.00)] 99.93 (0.03) [87.37 (0.66)] 99.93 (0.03) [88.08 (0.58)] 99.18 (0.11) [74.03 (4.05)] 99.95 (0.03) [83.21 (0.89)] 99.93 (0.01) [80.38 (1.60)] 99.85 (0.04) [80.02 (3.75)]
CIFAR-10 90.20 (0.29) [100.00 (0.00)] 97.22 (0.18) [80.27 (0.23)] 97.17 (0.17) [80.53 (0.21)] 91.40 (0.89) [74.42 (16.73)] 95.45 (0.95) [80.56 (1.14)] 93.72 (0.94) [80.57 (4.78)] 94.25 (0.31) [80.91 (0.83)]
OrganMNIST 89.10 (1.06) [100.00 (0.00)] 96.55 (0.78) [80.25 (1.42)] 96.52 (0.82) [80.33 (1.18)] 93.79 (0.95) [81.77 (6.86)] 94.49 (0.65) [80.17 (4.32)] 93.47 (0.55) [80.16 (4.59)] 93.68 (1.73) [80.04 (2.48)]
OctMNIST 91.93 (0.22) [100.00 (0.00)] 97.08 (0.20) [80.94 (0.28)] 97.18 (0.19) [80.22 (0.20)] 93.43 (0.81) [85.57 (8.22)] 95.40 (1.06) [81.05 (2.58)] 94.66 (0.31) [85.32 (4.92)] 94.91 (0.79) [86.67 (3.26)]
Noisy (25%) HAR 96.51 (0.44) [100.00 (0.00)] 98.56 (1.66) [81.16 (15.40)] 98.56 (1.66) [81.13 (15.41)] 97.22 (0.55) [80.24 (7.73)] 97.82 (0.35) [80.50 (0.67)] 97.78 (1.00) [69.11 (6.36)] 98.85 (0.43) [80.18 (3.11)]
Gas Drift 93.84 (1.81) [100.00 (0.00)] 97.30 (2.60) [80.02 (16.41)] 97.28 (2.59) [80.09 (16.38)] 95.87 (2.48) [81.51 (9.20)] 98.71 (0.27) [80.32 (1.91)] 99.02 (0.48) [76.54 (2.11)] 97.52 (0.75) [75.31 (2.76)]
MNIST 97.88 (0.23) [100.00 (0.00)] 99.89 (0.01) [80.44 (1.54)] 99.89 (0.01) [80.64 (1.36)] 98.00 (0.20) [92.95 (10.17)] 99.94 (0.03) [80.38 (1.08)] 99.93 (0.03) [81.45 (0.58)] 99.95 (0.03) [81.20 (1.55)]
CIFAR-10 85.31 (0.18) [100.00 (0.00)] 92.25 (0.33) [81.61 (0.98)] 92.50 (0.38) [80.71 (1.00)] 85.84 (1.17) [88.25 (23.07)] 89.58 (1.28) [81.83 (2.48)] 90.93 (0.45) [80.29 (1.59)] 92.22 (0.12) [80.55 (0.41)]
OrganMNIST 89.10 (0.77) [100.00 (0.00)] 95.86 (1.36) [82.07 (1.95)] 96.29 (0.94) [80.93 (1.16)] 93.40 (0.74) [83.78 (5.13)] 96.74 (0.68) [81.19 (0.77)] 94.67 (1.17) [80.39 (5.45)] 94.48 (0.77) [80.39 (3.51)]
OctMNIST 91.89 (0.20) [100.00 (0.00)] 97.17 (0.21) [80.29 (0.87)] 97.10 (0.21) [80.72 (0.78)] 93.42 (1.10) [83.14 (12.31)] 95.49 (1.35) [81.20 (4.24)] 94.08 (0.64) [89.10 (2.42)] 94.63 (1.14) [77.56 (13.79)]
Target Coverage: 90% Clean HAR 97.38 (0.34) [100.00 (0.00)] 99.56 (0.17) [90.34 (0.38)] 99.58 (0.17) [90.19 (0.42)] 98.19 (0.74) [90.20 (14.01)] 98.75 (0.73) [90.89 (1.26)] 99.23 (0.19) [90.34 (0.29)] 99.37 (0.24) [90.27 (0.75)]
Gas Drift 94.10 (1.87) [100.00 (0.00)] 98.05 (1.71) [90.24 (1.58)] 98.12 (1.58) [90.06 (1.37)] 96.64 (2.06) [90.30 (6.65)] 96.96 (0.56) [88.09 (0.70)] 97.46 (0.44) [84.51 (2.08)] 96.26 (0.53) [87.24 (2.24)]
MNIST 98.55 (0.19) [100.00 (0.00)] 99.89 (0.04) [90.57 (0.55)] 99.89 (0.04) [90.84 (0.53)] 98.87 (0.13) [91.82 (1.52)] 99.93 (0.02) [90.42 (0.36)] 99.80 (0.03) [90.22 (0.39)] 99.84 (0.03) [90.72 (3.41)]
CIFAR-10 90.20 (0.29) [100.00 (0.00)] 94.41 (0.21) [90.19 (0.19)] 94.43 (0.21) [90.14 (0.20)] 90.32 (0.22) [93.66 (10.05)] 92.77 (0.68) [90.16 (0.72)] 92.98 (0.86) [91.01 (1.63)] 91.54 (0.44) [90.43 (0.47)]
OrganMNIST 89.10 (1.06) [100.00 (0.00)] 92.60 (0.87) [90.15 (1.43)] 92.57 (0.94) [90.27 (1.20)] 91.44 (1.68) [90.49 (7.46)] 92.52 (0.66) [90.42 (1.78)] 92.35 (0.58) [90.16 (1.68)] 90.88 (0.75) [90.07 (2.00)]
OctMNIST 91.93 (0.22) [100.00 (0.00)] 94.98 (0.26) [91.35 (0.43)] 95.17 (0.25) [90.69 (0.38)] 92.23 (0.60) [96.63 (4.21)] 91.96 (1.94) [91.50 (1.62)] 93.49 (0.46) [91.18 (2.08)] 94.20 (0.68) [90.38 (2.12)]
Noisy (25%) HAR 96.51 (0.44) [100.00 (0.00)] 98.28 (1.44) [90.11 (8.10)] 98.29 (1.45) [90.40 (7.86)] 96.94 (0.77) [88.55 (4.82)] 98.48 (0.29) [84.94 (1.07)] 97.78 (1.00) [69.11 (6.36)] 98.56 (0.35) [89.34 (1.23)]
Gas Drift 93.84 (1.81) [100.00 (0.00)] 96.58 (2.05) [90.75 (7.96)] 96.68 (2.11) [90.04 (8.34)] 95.11 (1.85) [91.71 (6.72)] 98.71 (0.27) [80.32 (1.91)] 99.02 (0.48) [76.54 (2.11)] 97.52 (0.75) [75.31 (2.76)]
MNIST 97.88 (0.23) [100.00 (0.00)] 99.71 (0.03) [90.01 (0.99)] 99.70 (0.04) [90.03 (0.89)] 98.00 (0.20) [92.95 (10.17)] 99.82 (0.06) [90.33 (0.61)] 99.89 (0.03) [83.47 (0.43)] 99.86 (0.02) [90.81 (0.85)]
CIFAR-10 85.31 (0.18) [100.00 (0.00)] 89.43 (0.39) [90.44 (0.94)] 89.42 (0.39) [90.43 (0.92)] 85.66 (0.69) [91.06 (17.80)] 89.71 (1.01) [83.40 (1.44)] 89.94 (0.92) [83.03 (1.17)] 92.14 (0.24) [81.90 (1.23)]
OrganMNIST 89.10 (0.77) [100.00 (0.00)] 92.56 (0.62) [90.22 (1.70)] 92.53 (0.64) [90.29 (1.60)] 92.24 (0.76) [90.61 (2.35)] 91.92 (1.05) [90.22 (1.17)] 91.77 (0.94) [90.11 (2.06)] 92.17 (0.24) [90.91 (0.45)]
OctMNIST 91.89 (0.20) [100.00 (0.00)] 95.22 (0.21) [90.47 (0.52)] 95.23 (0.23) [90.45 (0.45)] 92.41 (0.36) [90.70 (8.09)] 93.28 (0.52) [91.49 (0.93)] 93.63 (0.40) [90.89 (3.05)] 93.39 (0.89) [90.22 (0.89)]